(Direct proof) A claim is given as a quantified statement: “The product of two odd numbers is an odd number” a) (1 point) Write the domain of the variables: b) (4 points) Write the statement using quantifiers and an implication of propositional functions: c) (15 points) Prove the statement by direct proof (Assume the hypothesis and derive the conclusion) 


1
Expert's answer
2021-06-21T12:50:01-0400

Let "P(x)" denotes the statement ""x" is an odd number", "Q(x,y)" denotes the statement ""x\cdot y" is an odd number".

a) It follows that the the domain of the variables is the set "\Z" of integers.


b) Let us write the statement using quantifiers and an implication of propositional functions: "\forall (x\in\Z)\forall(y\in\Z)\ (P(x)\land P(y)\to Q(x,y))"


c) Let us prove the statement by direct proof. Let "x\in \Z" and "y\in\Z" be odd integers. Then "x=2k+1,y=2t+1" for some "k,t\in\Z." It follows that "x\cdot y=(2k+1)\cdot(2t+1)=4kt+2k+2t+1=2(2kt+k+t)+1", and hence "x\cdot y" is an odd number.